OpenStudy (anonymous):

Yolanda has finished 45% of the homework problems she was assigned. So far she has finished 27 problems. A. Write 45% as a rate per 100 and then reduce the fraction to lowest terms. B. How many problems was Yolanda assigned? Justify your answer. C. How many problems does Yolanda still have to do? Show your work PLEASE HELP

OpenStudy (retireed):



OpenStudy (anonymous):

0.45

OpenStudy (anonymous):

@usercode3rror

OpenStudy (retireed):

No, What is the largest common multiple to both he numerator and the denominator? it is one of these... Is it 1, 2 , 3, 4, 5, 6, 7, 8, 9, or 10?

OpenStudy (anonymous):

i dont get this

OpenStudy (anonymous):

nvr mind

OpenStudy (retireed):

Okay

OpenStudy (shaik0124):

let assume x as the total number of questions now 45%(x) is completed which is given as 27 (45/100 ) *x=27 solve for x u will get the answer for q 2

OpenStudy (shaik0124):

@aliyahbufford1239442 do it now

OpenStudy (anonymous):

@shaik0124 i have other thing to do right now so uhh i dont have time to do this right now

OpenStudy (shaik0124):

45x=2700 x=2700/45 x=60 (answer for q 2) 60-27=33 (answer for q 3) @aliyahbufford1239442

OpenStudy (anonymous):

thanks